OpenStudy (anonymous):

Given the triangle below, which of the following is a correct statement? Right triangle ABC with AB measuring 4, AC measuring 5, and BC measuring 6. csc right triangle C right triangle equals fraction 3 over 2 sec right triangle C right triangle equals fraction 5 over 6 csc right triangle B right triangle equals fraction 3 over 2 sec right triangle B right triangle equals fraction 6 over 5

OpenStudy (anonymous):

1st statement seems to be correct.

OpenStudy (anonymous):

Yes 1st one is correct. csc of C would be \[\frac{ hyp }{ opp }\] so\[\frac{ 6 }{ 4 }\] which reduces to \[\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

Alright guys! Thank you so much! I'm really lost when it comes to math!